Here is the Problem:
I have a family of 4 (2 adults, 2 children). I need to compare the cost for between a daily pass and yearly pass for the aqaurium. The family will generally visit the aqauriom 5 times per year. Which option will be most advantageous based on the number of yearly visits.
The cost of the Yearly pass is as follows:
Adult Yearly pass is $62.95
Child yearly Pass is $45.95
The Daily pass is as follows:
Adult pass is $24.95
The child is $18.95
I need to develop a linear algebraic equation using either substituion or elimination method showing where the two the two cost options are equivilalent, showing all work.
I also then need to depict this on a graph.
Please help.
I can be reached at oasisrm@hotmail.com
Thanks Answer by stanbon(75887) 
 I have a family of 4 (2 adults, 2 children). I need to compare the cost for between a daily pass and yearly pass for the aquarium. The family will generally visit the aqaurium 5 times per year. Which option will be most advantageous based on the number of yearly visits.
The cost of the Yearly pass is as follows:
Adult Yearly pass is $62.95
Child yearly Pass is $45.95
---
Yearly pass cost: 2*62.95+2*45.95 = 217.80
--
The Daily pass is as follows:
Adult pass is $24.95
The child is $18.95
---------------------------------------
Let x be the number of days the family attends the aquarium.
Daily pass cost: x(2*24.95+2*18.95) = x(2*24.95+2*18.95) = 87.8x
----------------------------------------------------------------
I need to develop a linear algebraic equation using either substituion or elimination method showing where the two the two cost options are equivalent, showing all work.
I also then need to depict this on a graph.
------
Equations:
Yearly cost = 217.8
Daily cost = 87.8x
-----
Equivalent means the costs are the same:
Solve for "x":
87.8x=217.8
x = 2.48 days.
---
If the family goes to the aguarium 1 or 2 days the daily
pass is their best buy.
If the family goes to the aquarium 3,4, or 5 days the
yearly pass is the best buy.
================================
Cheers,
Stan H.